### v1

# TDSM: Triplet Diffusion for Skeleton-Text Matching in Zero-Shot Action Recognition

Abstract
We firstly present a diffusion-based action recognition with zero-shot learning for skeleton inputs. In zero-shot skeleton-based action recognition, aligning skeleton features with the text features of action labels is essential for accurately predicting unseen actions. Previous methods focus on direct alignment between skeleton and text latent spaces, but the modality gaps between these spaces hinder robust generalization learning. Motivated from the remarkable performance of text-to-image diffusion models, we leverage their alignment capabilities between different modalities mostly by focusing on the training process during reverse diffusion rather than using their generative power. Based on this, our framework is designed as a Triplet Diffusion for Skeleton-Text Matching (TDSM) method which aligns skeleton features with text prompts through reverse diffusion, embedding the prompts into the unified skeleton-text latent space to achieve robust matching. To enhance discriminative power, we introduce a novel triplet diffusion (TD) loss that encourages our TDSM to correct skeleton-text matches while pushing apart incorrect ones. Our TDSM significantly outperforms the very recent state-of-the-art methods with large margins of 2.36%-point to 13.05%-point, demonstrating superior accuracy and scalability in zero-shot settings through effective skeleton-text matching.
1 Introduction
Human action recognition [51, 43, 56, 25] focuses on classifying actions from movements, with RGB videos commonly used due to their accessibility. However, recent advancements in depth sensors [22] and pose estimation algorithms [4, 50] have driven the adoption of skeleton-based action recognition. Skeleton data offers several advantages: it captures only human poses without background noise, ensuring a compact representation. Furthermore, 3D skeletons remain invariant to environmental factors such as lighting, background, and camera angles, providing consistent 3D coordinates across conditions [12].
Figure 1: Motivation of the proposed TDSM pipeline. Previous methods rely on direct alignment between skeleton and text latent spaces, but suffer from modality gaps that limit generalization. Our TDSM overcomes this challenge by leveraging a reverse diffusion process to embed text prompts into the unified skeleton-text latent space, ensuring more effective cross-modal alignment.
Despite these benefits, scaling skeleton-based recognition to large action sets remains challenging. Traditional methods [63, 8, 12, 70, 11, 67, 6, 9] require training on annotated data for each action, which is impractical for a large number of possible actions. Zero-shot skeleton-based action recognition [47, 16, 69, 71, 32, 7, 61, 26, 30] addresses this by enabling predictions for unseen actions without requiring explicit training data, making it valuable for applications such as surveillance, robotics, and human-computer interaction, where continuous learning is infeasible [57, 15].
Zero-shot models offer the flexibility for unseen data by aligning pre-learned skeleton features with text-based descriptions corresponding to action labels, ensuring their scalability in real-world scenarios. However, achieving the effective alignment between skeleton data and text features entails significant challenges. While skeleton data captures temporal and spatial motion patterns, the text descriptions for action labels carry high-level semantic information. This modality gap makes it difficult to align their corresponding latent spaces effectively, hindering the generalization learning for unseen actions—particularly when distinguishing between semantically similar actions (e.g., “Throw” vs. “Shoot”).
Fig. 1 illustrates the key differences between previous methods and our proposed method. The previous methods [47, 16, 69, 71, 32, 7, 61, 26, 30] aim at directly aligning the different features of skeletons and text descriptions for action labels between the skeleton and text latent spaces, which struggles to generalize due to the inherent differences between these modalities in different latent spaces. On the other hand, our TDSM leverages a reverse diffusion training scheme to merge the skeleton features with their corresponding text prompts, which leads to discriminatively fused features in a unified latent space.
Inspired by text-to-image diffusion models [45, 14], we propose a novel framework: Triplet Diffusion for Skeleton-Prompt Matching (TDSM). Unlike to be a generative model, our TDSM leverages the reverse diffusion training scheme to merge skeleton features with text prompts which aims at achieving merged discriminative (not generative) features for improved action recognition tasks. More specifically, our TDSM learns to denoise noisy skeleton features conditioned on the corresponding text prompts, embedding the prompts into a merged skeleton-text latent space to ensure effective implicit alignment. This approach avoids the challenges of direct latent space alignment while enhancing robustness. Additionally, we introduce a novel triplet diffusion (TD) loss, which encourages tighter alignment for correct skeleton-text pairs and pushes apart incorrect ones, further improving the model’s discriminative power. As an additional benefit, the diffusion process, driven by the random noise added during training, acts as a natural regularization mechanism. This prevents overfitting and enhances the model’s ability to generalize effectively to unseen actions. Our contributions are threefold:
  • We firstly present a diffusion-based action recognition with zero-shot learning for skeleton inputs, called a Triplet Diffusion for Skeleton-Text Matching (TDSM) which is the first framework to apply diffusion models and to implicitly align the skeleton features with text prompts by fully taking the advantage of excellent text-image correspondence learning in generative diffusion process, thus being able to learn fused discriminative features in a unified latent space.
  • We propose a novel triplet diffusion (TD) loss to enhance the model’s discriminative power by ensuring accurate denoising for correct skeleton-text pairs while suppressing incorrect pairs.
  • Our TDSM significantly outperforms the very recent state-of-the-art (SOTA) methods with large margins of 2.36%-point to 13.05%-point across multiple benchmarks, demonstrating scalability and robustness under various seen-unseen split settings.
2 Related Work
2.1 Skeleton-based Action Recognition
Traditional skeleton-based action recognition assumes fully annotated training and test datasets, in contrast to other skeleton-based action recognition methods under zero-shot settings which aim to recognize unseen classes without explicit training samples. Early methods [66, 34, 72, 31] employed RNN-based models to capture the temporal dynamics of skeleton sequences. Subsequent studies [3, 62, 23, 12] explored CNN-based approaches, transforming skeleton data into pseudo-images. Recent advancements leverage graph convolutional networks (GCNs) [6, 67, 28, 9, 60, 36, 27, 70] to effectively represent the graph structures of skeletons, comprising joints and bones. ST-GCN [63] introduced graph convolutions along the skeletal axis combined with 1D temporal convolutions to capture motion over time. Shift-GCN [8] improved computational efficiency by implementing shift graph convolutions. Building on these methods, transformer-based models [68, 39, 13, 55, 11] have been proposed to address the limited receptive field of GCNs by capturing global skeletal-temporal dependencies. In this work, we adopt ST-GCN [63] and Shift-GCN [8] to extract skeletal-temporal representations from skeleton data, transforming input skeleton sequences into a latent space for further processing in the proposed framework.
2.2 Zero-shot Skeleton-based Action Recognition
Most of the existing works focus on aligning the skeleton latent space with the text latent space. These approaches can be categorized broadly into VAE-based methods [47, 16, 30, 32] and contrastive learning-based methods [69, 71, 7, 26, 61].
VAE-based. The previous work, CADA-VAE [47], leverages VAEs [24] to align skeleton and text latent spaces, ensuring that each modality’s decoder can generate useful outputs from the other’s latent representation. SynSE [16] refines this by introducing separate VAEs for verbs and nouns, improving the structure of the text latent space. MSF [30] extends this approach by incorporating action and motion-level descriptions to enhance alignment. SA-DVAE [32] disentangles skeleton features into semantic-relevant and irrelevant components, aligning text features exclusively with relevant skeleton features for improved performance.
Contrastive learning-based. Contrastive learning-based methods align skeleton and text features through positive and negative pairs [5]. SMIE [69] concatenates skeleton and text features, and applies contrastive learning by treating masked skeleton features as positive samples and other actions as negatives. PURLS [71] incorporates GPT-3 [1] to generate text descriptions based on body parts and motion evolution, using cross-attention to align text descriptions with skeleton features. STAR [7] extends this idea with GPT-3.5 [1], generating text descriptions for six distinct skeleton groups, and introduces learnable prompts to enhance alignment. DVTA [26] introduces a dual alignment strategy, performing direct alignment between skeleton and text features, while also generating augmented text features via cross-attention for improved alignment. InfoCPL [61] strengthens contrastive learning by generating 100 unique sentences per action label, enriching the alignment space.
While most existing methods rely on direct alignment between skeleton and text latent spaces, they often struggle with generalization due to inherent differences between the two modalities. To address these limitations, our TDSM leverages a novel triplet diffusion approach. It uses text prompts to guide the reverse diffusion process, embedding these prompts into the unified skeleton-text latent space for more effective implicit alignment via feature fusion. Also, our triplet diffusion loss, inspired by the triplet loss [19], further enhances the model’s discriminative power by promoting correct alignments and penalizing incorrect ones.
Figure 2: Training framework of our TDSM for zero-shot skeleton-based action recognition.
2.3 Diffusion Models
Diffusion models have become a fundamental to generative tasks by learning to reverse a noise-adding process for original data recovery. Denoising Diffusion Probabilistic Models (DDPMs) [18] introduced a step-by-step denoising framework, enabling the modeling of complex data distributions and establishing the foundation for diffusion-based generative models. Building on this, Latent Diffusion Models (LDMs) [45, 14] improve computational efficiency by operating in lower-dimensional latent spaces while maintaining high-quality outputs. LDMs have been particularly successful in text-to-image (T2I) generation, showing the potential of diffusion models for cross-modal alignment tasks. In the reverse diffusion process, LDMs employ a denoising U-Net [46] where text prompts are integrated with image features through cross-attention blocks, effectively guiding the model to align the two modalities. Further extending this line of research, Diffusion Transformers (DiTs) [41] integrate transformer architectures into diffusion processes, capturing long-range dependencies across spatial dimension. In this work, we leverage the aligned fusion capabilities of diffusion models, focusing on the learning process during reverse diffusion rather than their generative power. Specifically, we utilize a DiT-based network as a denoising model, where text prompts guide the denoising of noisy skeleton features. This approach embeds text prompts into the unified latent space in the reverse diffusion process, ensuring robust fusion of the two modalities and enabling effective generalization to unseen actions in zero-shot recognition settings.
Zero-shot tasks with diffusion models. Recently, diffusion models have also been extended to zero-shot tasks in RGB-based vision applications, such as semantic correspondence [65], segmentation [52, 2], image captioning [64], and image classification [10, 29]. These approaches often rely on large-scale pretrained diffusion models, such as LDMs [45, 14], trained on datasets like LAION-5B [48] with billions of text-image pairs. In contrast, our approach demonstrates that diffusion models can be effectively applied to smaller, domain-specific tasks, such as skeleton-based action recognition, without the need for large-scale finetuning. This highlights the versatility of diffusion models beyond large-scale vision-language tasks, providing a practical solution to zero-shot generalization with limited data resources.
3 Methods
3.1 Overview of TDSM
In the training phase, we are given a dataset
$\mathcal{D}_{\text{train}}=\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{N},\quad y_{i}\in\mathcal{Y},$ (1)
where $\mathbf{X}_{i}\in\mathbb{R}^{T\times V\times M\times C_{\text{in}}}$ represents a skeleton sequence, and $y_{i}$ is the corresponding ground truth label. Each skeleton sequence $\mathbf{X}_{i}$ consists of sequence length $T$, the number $V$ of joints, the number $M$ of actors, and the dimensionality $C_{\text{in}}$ representing each joint. The label $y_{i}$ belongs to the set of seen class labels $\mathcal{Y}$. Here, $N$ denotes the total number of training samples in the seen dataset. In the inference phase, we are provided with a test dataset
$\mathcal{D}_{\text{test}}=\{(\mathbf{X}_{j}^{u},y_{j}^{u})\}_{j=1}^{N_{u}},\quad y_{j}^{u}\in\mathcal{Y}_{u},$ (2)
where $\mathbf{X}_{j}^{u}\in\mathbb{R}^{T\times V\times M\times C_{\text{in}}}$ denotes skeleton sequences from unseen classes, and $y_{j}^{u}$ are their corresponding labels. In this phase, $N_{u}$ represents the total number of test samples from unseen classes. In the zero-shot setting, the seen and unseen label sets are disjoint, i.e.,
$\mathcal{Y}\cap\mathcal{Y}_{u}=\varnothing.$ (3)
We train the TDSM using $\mathcal{D}_{\text{train}}$ and enable it to generalize to unseen classes from $\mathcal{D}_{\text{test}}$. By learning a robust discriminative fusion of skeleton features and text descriptions, the model can predict the correct label ${\hat{y}}^{u}\in\mathcal{Y}_{u}$ for an unseen skeleton sequence $\mathbf{X}_{j}^{u}$ during inference.
Fig. 2 provides an overview of our training framework of TDSM. The framework consists of three main components: (i) skeleton and text encoders, (ii) the forward process, and (iii) the reverse process. As detailed in Sec. 3.2, the pretrained skeleton encoder $\mathcal{E}_{x}$ and text encoder $\mathcal{E}_{d}$ embed the skeleton inputs $\mathbf{X}$ and prompt input $\mathbf{d}$ with an action label $y$ into their respective feature spaces, producing the skeleton feature $\mathbf{z}_{x}$ and two types of text features: the global text feature $\mathbf{z}_{g}$ and the local text feature $\mathbf{z}_{l}$. The skeleton feature $\mathbf{z}_{x}$ undergoes the forward process, where noise $\bm{\epsilon}$ is added to it. In the reverse process, as described in Sec. 3.3, the Diffusion Transformer $\mathcal{T}_{\text{diff}}$ predicts the noise $\hat{\bm{\epsilon}}$. The training objective function that ensures the TDSM to learn robust discriminative power is discussed in Sec. 3.3. Finally, Sec. 3.4 explains the strategy used during the inference phase to predict the correct label ${\hat{y}}^{u}$ for unseen actions.
3.2 Embedding Skeleton and Prompt Input
Following the LDMs [45, 14], we perform the diffusion process in a compact latent space by projecting both skeleton data and prompt into their respective feature spaces. For skeleton data, we adopt GCNs, specifically ST-GCN [63] and Shift-GCN [8], as the architecture for the skeleton encoder $\mathcal{E}_{x}$ that is trained on $\mathcal{D}_{\text{train}}$ using a cross-entropy loss:
$\mathcal{L}_{\text{CE}}=-\sum_{k=1}^{|\mathcal{Y}|}\mathbf{y}(k)\log\hat{\mathbf{y}}(k),$ (4)
where $\hat{\mathbf{y}}=\mathsf{MLP}(\mathcal{E}_{x}(\mathbf{X}))$ is the predicted class label for a skeleton input $\mathbf{X}_{i}$, $|\mathcal{Y}|$ is the number of seen classes and $\mathbf{y}$ is the one-hot vector of the ground truth label $y$. Once trained, the parameters of skeleton encoder $\mathcal{E}_{x}$ are frozen and used to generate the skeleton latent space representation $\mathbf{z}_{x}=\mathcal{E}_{x}(\mathbf{X})$. After reshaping the feature for the attention layer, $\mathbf{z}_{x}$ is represented in $\mathbb{R}^{M_{x}\times C}$, where $M_{x}$ is the number of skeleton tokens, and $C$ is the feature dimension.
For text encoder $\mathcal{E}_{d}$, we leverage the text prompts provided by SA-DVAE [32] to capture rich semantic information about the action labels. Each ground truth (GT) label is associated with a prompt $\mathbf{d}_{p}$, while a wrong label (negative sample) is assigned a prompt $\mathbf{d}_{n}$. To encode these prompts, we utilize a pretrained text encoder, such as CLIP [44, 21], which provides two types of output features: a global text feature $\mathbf{z}_{g}$ and a local text feature $\mathbf{z}_{l}$. The text encoder’s output for a given prompt d can be represented as:
$\left[\mathbf{z}_{g}\mid\mathbf{z}_{l}\right]=\mathcal{E}_{d}(\textbf{d}),$ (5)
[…]
For each GT label (positive sample), the text encoder $\mathcal{E}_{d}$ extracts both the global and local text features, denoted as $\mathbf{z}_{g,p}$ and $\mathbf{z}_{l,p}$, respectively. Similarly, for each wrong label (negative sample), the encoder extracts the features $\mathbf{z}_{g,n}$ and $\mathbf{z}_{l,n}$. These four features later guide the diffusion process by conditioning the denoising of noisy skeleton features.
3.3 Diffusion Process
Our framework leverages a conditional denoising diffusion process, not to generate data but to learn a discriminative skeleton latent space by fusing skeleton features with text prompts through the reverse diffusion process. Our TDSM is trained to denoise skeleton features such that the resulting latent space becomes discriminative with respect to action labels. Guided by our triplet diffusion (TD) loss, the denoising process conditions on text prompts to strengthen the discriminative fusion of skeleton features and their corresponding prompts. The TD loss encourages correct skeleton-text pairs to be pulled closer in the fused skeleton-text latent space while pushing apart incorrect pairs, enhancing the model’s discriminative power.
Forward process. Random Gaussian noise is added to the skeleton feature $\mathbf{z}_{x}$ at a random timestep $t\sim\mathcal{U}(T)$ within total $T$ steps. At each randomly selected step $t$, the noisy feature $\mathbf{z}_{x,t}$ is generated as:
$\mathbf{z}_{x,t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{x}+\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon},$ (6)
where $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is Gaussian noise, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ controls the noise level at step $t$ [18].
Reverse process. The Diffusion Transformer $\mathcal{T}_{\text{diff}}$ predicts noise $\hat{\bm{\epsilon}}$ from noisy feature $\mathbf{z}_{x,t}$, conditioned on the global and local text features $\mathbf{z}_{g}$ and $\mathbf{z}_{l}$ at given timestep $t$:
$\hat{\bm{\epsilon}}=\mathcal{T}_{\text{diff}}\left(\mathbf{z}_{x,t},\mathbf{z}_{g},\mathbf{z}_{l},t\right).$ (7)
Using the shared weights in $\mathcal{T}_{\text{diff}}$, we predict $\hat{\bm{\epsilon}}_{p}$ for positive features $\left(\mathbf{z}_{g,p},\mathbf{z}_{l,p}\right)$ and $\hat{\bm{\epsilon}}_{n}$ for negative features $\left(\mathbf{z}_{g,n},\mathbf{z}_{l,n}\right)$.
Diffusion Transformer $\mathcal{T}_{\text{diff}}$. The Diffusion Transformer $\mathcal{T}_{\text{diff}}$ takes $\mathbf{z}_{x,t}$, $\mathbf{z}_{g}$, $\mathbf{z}_{l}$, and $t$ as inputs (Fig. 2). These inputs are embedded into corresponding feature representations $\mathbf{f}_{x,t}$, $\mathbf{f}_{c}$, and $\mathbf{f}_{l}$ as follows:
$\begin{split}&\mathbf{f}_{x,t}=\mathsf{Linear}(\mathbf{z}_{x,t})+\mathsf{PE}_{x},\\ &\mathbf{f}_{c}=\mathsf{Linear}(\mathsf{TE}_{t})+\mathsf{Linear}(\mathbf{z}_{g}),\\ &\mathbf{f}_{l}=\mathsf{Linear}(\mathbf{z}_{l})+\mathsf{PE}_{l},\\ \end{split}$ (8)
where $\mathsf{PE}_{x}$ and $\mathsf{PE}_{l}$ are positional embeddings applied to the feature maps, capturing spatial positional information, while $\mathsf{TE}_{t}$ is a timestep embedding [53] that maps the scalar $t$ to a higher-dimensional space. The embedded features $\mathbf{f}_{x,t}$, $\mathbf{f}_{c}$, and $\mathbf{f}_{l}$ are then passed through $B$ CrossDiT Blocks, followed by a Layer Normalization ($\mathsf{LN}$) and a final $\mathsf{Linear}$ layer to predict the noise $\hat{\bm{\epsilon}}\in\mathbb{R}^{M_{x}\times C}$.
CrossDiT Block. The CrossDiT Block facilitates interaction between skeleton and text features, enhancing fusion through effective feature modulation [58, 42] and multi-head self-attention [54]. Fig. 3 shows a detail structure of our CrossDiT Block. Built upon the DiTs architecture [41, 14], it leverages modulation techniques and self-attention mechanisms to efficiently capture the dependencies across these modalities. The skeleton feature $\mathbf{f}_{x}$ and local text feature $\mathbf{f}_{l}$ are first modulated separately using $\mathsf{Scale}$-$\mathsf{Shift}$ and $\mathsf{Scale}$ operations as:
$\begin{split}&\left[\bm{\alpha}_{x}\mid\bm{\beta}_{x}\mid\bm{\gamma}_{x}\>|\>\bm{\alpha}_{l}\mid\bm{\beta}_{l}\mid\bm{\gamma}_{l}\right]=\mathsf{Linear}(\mathbf{f}_{c}),\\ &\mathsf{Scale}\text{-}\mathsf{Shift}:\;\mathbf{f}_{i}\leftarrow(1+\bm{\gamma}_{i})\odot\mathbf{f}_{i}+\bm{\beta}_{i},\\ &\mathsf{Scale}:\;\mathbf{f}_{i}\leftarrow\bm{\alpha}_{i}\odot\mathbf{f}_{i},\\ \end{split}$ (9)
where $i\in\left\{x,l\right\}$ denotes the skeleton or local text feature, respectively. The parameters $\bm{\alpha}$, $\bm{\beta}$, and $\bm{\gamma}$ are conditioned on the global text feature $\mathbf{z}_{g}$ and timestep $t$, allowing the block to modulate feature representations effectively. Also, we compute query, key, and value matrices for both skeleton and local text features separately:
$[\mathbf{q}_{i}\mid\mathbf{k}_{i}\mid\mathbf{v}_{i}]=\mathsf{Linear}(\mathbf{f}_{i}).$ (10)
These matrices are token-wise concatenated and fed into a multi-head self-attention module, followed by a split to retain token-specific information as:
$[\mathbf{f}_{x}\mid\mathbf{f}_{l}]\leftarrow\mathsf{SoftMax}\left(\left[\mathbf{q}_{x}\mid\mathbf{q}_{l}\right]\left[\mathbf{k}_{x}\mid\mathbf{k}_{l}\right]^{\mathsf{T}}\right)\left[\mathbf{v}_{x}\mid\mathbf{v}_{l}\right].$ (11)
By leveraging the attention from skeleton, timestep, and text features, the CrossDiT Block ensures efficient interaction between modalities, promoting the skeleton-text fusion for discriminative feature learning and improved generalization to unseen actions.
Figure 3: A detail structure of our CrossDiT Block.
Loss function design. The overall training objective combines a diffusion loss and our triplet diffusion (TD) loss to promote both effective noise prediction and discriminative alignment (fusion). The total loss is defined as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{diff}}+\lambda\mathcal{L}_{\text{TD}},$ (12)
where $\mathcal{L}_{\text{diff}}$ ensures accurate denoising, and $\mathcal{L}_{\text{TD}}$ enhances the ability to differentiate between correct and incorrect label predictions. The diffusion loss $\mathcal{L}_{\text{diff}}$ [18] is given by:
$\mathcal{L}_{\text{diff}}=\|\bm{\epsilon}-\hat{\bm{\epsilon}}_{p}\|_{2},$ (13)
where $\bm{\epsilon}$ is a true noise, and $\bm{\hat{\epsilon}}_{p}$ is a predicted noise for the GT text feature. Our triplet diffusion loss $\mathcal{L}_{\text{TD}}$ is defined as:
$\mathcal{L}_{\text{TD}}=\max\left(\|\bm{\epsilon}-\hat{\bm{\epsilon}}_{p}\|_{2}-\|\bm{\epsilon}-\hat{\bm{\epsilon}}_{n}\|_{2}+\tau,\;0\right),$ (14)
where $\hat{\bm{\epsilon}}_{n}$ is the predicted noise for an incorrect (negative) text feature, and $\tau$ is a margin parameter. $\mathcal{L}_{\text{TD}}$ is simple but very effective, encouraging the model to minimize the distance ($\|\bm{\epsilon}-\hat{\bm{\epsilon}}_{p}\|_{2}$) between true noise $\bm{\epsilon}$ and GT prediction $\hat{\bm{\epsilon}}_{p}$ while maximizing the distance ($\|\bm{\epsilon}-\hat{\bm{\epsilon}}_{n}\|_{2}$) for negative predictions $\hat{\bm{\epsilon}}_{n}$, which can ensure discriminative fusion of two modalities in the learned skeleton-text latent space.
Figure 4: Inference framework of our TDSM for zero-shot skeleton-based action recognition.
3.4 Inference Phase
Our approach enhances discriminative fusion through the TD loss, which is designed to denoise GT skeleton-text pairs effectively while preventing the fusion of incorrect pairs within the seen dataset. This selective denoising process promotes a robust fusion of skeleton and text features, allowing the model to develop a discriminative feature space that can generalize to unseen action labels.
Importantly, human actions often share common skeletal movements across related actions, and our model leverages them within seen skeleton data to extrapolate and associate unseen skeleton sequences with their corresponding text prompts. This capability to infer from familiar patterns reinforces the model’s discriminative power, enabling reliable zero-shot recognition of previously unseen actions.
In inference, each unseen skeleton sequence $\mathbf{X}^{u}$ and its all candidate text prompts are inputted to our TDSM, and the resulting noises for the all candidate text prompts are compared with a fixed GT noise $\bm{\epsilon}_{\text{test}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. $\mathbf{X}^{u}$ is first encoded into the skeleton latent space through $\mathcal{E}_{x}$ as:
$\mathbf{z}_{x}^{u}=\mathcal{E}_{x}(\mathbf{X}^{u}).$ (15)
Each candidate action label $y_{k}^{u}\in\mathcal{Y}_{u}$ is associated with a prompt $\mathbf{d}_{k}^{u}$ that is processed through $\mathcal{E}_{d}$ to extract the global and local text features:
$\left[\mathbf{z}_{g,k}^{u}\mid\mathbf{z}_{l,k}^{u}\right]=\mathcal{E}_{d}(\mathbf{d}_{k}^{u}).$ (16)
Next, the forward process is performed using a fixed Gaussian noise $\bm{\epsilon}_{\text{test}}$ and a fixed timestep $t_{\text{test}}$ to generate the noisy skeleton feature as:
$\mathbf{z}_{x,t}^{u}=\sqrt{\bar{\alpha}_{t_{\text{test}}}}\,\mathbf{z}_{x}^{u}+\sqrt{1-\bar{\alpha}_{t_{\text{test}}}}\,\bm{\epsilon}_{\text{test}}.$ (17)
For $y_{k}^{u}$, the Diffusion Transformer $\mathcal{T}_{\text{diff}}$ predicts noise $\hat{\bm{\epsilon}}_{k}$ as:
$\hat{\bm{\epsilon}}_{k}=\mathcal{T}_{\text{diff}}(\mathbf{z}_{x,t}^{u},\mathbf{z}_{g,k}^{u},\mathbf{z}_{l,k}^{u},t_{\text{test}}).$ (18)
The score for $y_{k}^{u}$ is computed as the $\ell_{2}$-norm between $\bm{\epsilon}_{\text{test}}$ and $\hat{\bm{\epsilon}}_{k}$. The predicted label ${\hat{y}}^{u}$ is then the one that minimizes this distance:
${\hat{y}}^{u}=\arg\min_{k}\|\bm{\epsilon}_{\text{test}}-\hat{\bm{\epsilon}}_{k}\|_{2}.$ (19)
This process ensures that the model selects the action label whose text prompt well aligns with the skeleton sequence, enabling accurate zero-shot action recognition for unseen skeleton sequences with unseen action labels.
4 Experiments
4.1 Datasets
NTU RGB+D [49]. NTU RGB+D dataset, referred to as NTU-60, is one of the largest benchmarks for human action recognition, consisting of 56,880 action samples across 60 action classes. It captures 3D skeleton data, depth maps, and RGB videos using Kinect sensors [22], making it a standard for evaluating single- and multi-view action recognition models. The dataset provides a cross-subject split (X-sub) with 40 subjects in total, where 20 subjects are used for training and the remaining 20 for testing. In our experiments, we construct $\mathcal{D}_{\text{train}}$ from the X-sub training set with seen labels and $\mathcal{D}_{\text{test}}$ from the X-sub test set with unseen labels, ensuring a fully zero-shot action recognition setting.
NTU RGB+D 120 [35]. NTU RGB+D 120 dataset, referred to as NTU-120, extends the original NTU-60 dataset by adding 60 additional action classes, resulting in a total of 120 classes and 114,480 video samples. The X-sub setting for NTU-120 includes 106 subjects, with 53 used for training and the remaining 53 for testing. We apply the same experimental protocol as NTU-60, using the training set to form $\mathcal{D}_{\text{train}}$ with seen labels and the test set to construct $\mathcal{D}_{\text{test}}$ with unseen labels.
PKU-MMD [33]. PKU-MMD dataset is a large-scale dataset designed for multi-modality action recognition, offering 3D skeleton data and RGB+D recordings. The dataset contains a total of 66 subjects, where 57 subjects are used for training and the remaining 9 for testing. We follow the cross-subject setting to evaluate the generalization of our framework, constructing $\mathcal{D}_{\text{train}}$ with seen labels and $\mathcal{D}_{\text{test}}$ with unseen labels.
  Methods Publications NTU-60 (Acc, %) NTU-120  (Acc, %)
55/5 split 48/12 split 40/20 split 30/30 split 110/10 split 96/24 split 80/40 split 60/60 split
ReViSE [20] ICCV 2017 53.91 17.49 24.26 14.81 55.04 32.38 19.47 8.27
JPoSE [59] ICCV 2019 64.82 28.75 20.05 12.39 51.93 32.44 13.71 7.65
CADA-VAE [47] CVPR 2019 76.84 28.96 16.21 11.51 59.53 35.77 10.55 5.67
SynSE [16] ICIP 2021 75.81 33.30 19.85 12.00 62.69 38.70 13.64 7.73
SMIE [69] ACM MM 2023 77.98 40.18 - - 65.74 45.30 - -
PURLS [71] CVPR 2024 79.23 40.99 31.05 23.52 71.95 52.01 28.38 19.63
SA-DVAE [32] ECCV 2024 82.37 41.38 - - 68.77 46.12 - -
STAR [7] ACM MM 2024 81.40 45.10 - - 63.30 44.30 - -
TDSM (Ours) - 86.49 56.03 36.09 25.88 74.15 65.06 36.95 27.21
Table 1: Top-1 accuracy results of various zero-shot skeleton-based action recognition methods evaluated on the SynSE and PURLS benchmarks for the NTU-60 and NTU-120 datasets. Each split is denoted as X/Y, where X represents the number of seen classes and Y the number of unseen classes. The results in red highlight the best-performing model, while those in blue indicate the second-best. For our TDSM framework, the reported accuracy is the average value obtained from 10 trials, each with different Gaussian noise.
4.2 Experiment Setup
Our TDSM were implemented in PyTorch [40] and conducted on a single NVIDIA GeForce RTX 3090 GPU. Our model variants were trained for 50,000 iterations, with a warm-up period of 100 steps. We employed the AdamW optimizer [38] with a learning rate of $1\times 10^{-4}$ and a weight decay of 0.01. A cosine-annealing scheduler [37] was used to dynamically update the learning rate at each iteration. The batch size was set to 256, but for the TD loss computation, the batch was duplicated for positive and negative samples, resulting in an effective batch size of 512. Through empirical validation, the loss weight $\lambda$ and margin $\tau$ were both set to 1.0. The diffusion process was trained with a total timestep of $T=50$. For inference, the optimal timestep $t_{\text{test}}=25$ was selected based on accuracy trends across datasets (Fig. 5). We utilized $B=12$ CrossDiT Blocks, each containing a multi-head self-attention module with 12 heads. All feature dimensions were set to $C=768$. The local text features contained $M_{l}=35$ tokens, while the skeleton features were represented by a single token $M_{x}=1$. To ensure reproducibility, the random seed was fixed at 2,025 throughout all experiments. For the SynSE [16] and PURLS [71] seen and unseen split settings, we used the Shift-GCN [8] architecture as our skeleton encoder. In the SMIE [69] split setting, we employed the ST-GCN [63] structure. Additionally, for the SynSE and SMIE settings, we utilized both the text prompts and skeleton features provided by SA-DVAE [32]. However, as the PURLS setting did not offer pre-trained skeleton features, we directly trained them using cross-entropy loss $\mathcal{L}_{\text{CE}}$. Across all tasks, we adopted the text encoder from CLIP-ViT-H/14 [44, 21] to transform text prompts into latent representations.
4.3 Performance Evaluation
Evaluation on SynSE [16] and PURLS [71] benchmarks. Table 1 presents the performance comparison on the SynSE and PURLS benchmark splits across the NTU-60 and NTU-120 datasets. The SynSE benchmark focuses on standard settings, offering 55/5 and 48/12 splits on NTU-60, and 110/10 and 96/24 splits on NTU-120. These settings assess the model’s ability to generalize across typical seen-unseen splits. In contrast, the PURLS benchmark presents more extreme cases with 40/20 and 30/30 splits on NTU-60, and 80/40 and 60/60 splits on NTU-120, introducing higher levels of complexity by increasing the proportion of unseen labels in the test set. To account for the stochastic nature of noise during inference, we averaged the 10 runs with different Gaussian noise realizations. As shown in Table 1, our TDSM significantly outperforms the very recent state-of-the-art results across all benchmark splits, demonstrating superior generalization and robustness for various splits. Specifically, TDSM outperforms the existing methods on both standard (SynSE) and extreme (PURLS) settings, with 4.12%-point, 9.93%-point, 5.04%-point, and 2.36%-point improvements in top-1 accuracy on the NTU-60 55/5, 48/12, 40/20, and 30/30 splits, respectively, compared to the second best models. Also compared to the second best models, our TDSM attains 2.20%-point, 13.05%-point, 8.57%-point, and 7.58%-point accuracy improvements on the NTU-120 splits, further validating its scalability to larger datasets and more complex unseen classes.
  Methods NTU-60 (Acc, %) NTU-120 (Acc, %) PKU-MMD (Acc, %)
55/5 split 110/10 split 46/5 split
ReViSE [20] 60.94 44.90 59.34
JPoSE [59] 59.44 46.69 57.17
CADA-VAE [47] 61.84 45.15 60.74
SynSE [16] 64.19 47.28 53.85
SMIE [69] 65.08 46.40 60.83
SA-DVAE [32] 84.20 50.67 66.54
STAR [7] 77.50 - 70.60
TDSM (Ours) 88.88 69.47 70.76
Table 2: Top-1 accuracy results of various zero-shot skeleton-based action recognition methods evaluated on the NTU-60, NTU-120, and PKU-MMD datasets under the SMIE benchmark. The reported values are the average performance across three splits.
Evaluation on SMIE [69] benchmark. The SMIE benchmark provides three distinct splits to evaluate the generalization capability of models across different sets of unseen labels. Each split ensures that unseen labels do not overlap with seen ones, thereby rigorously testing the model’s ability to recognize new classes without prior exposure. For fair comparison, the reported performance is the average of the three splits. In this benchmark, as shown in Table 2, our TDSM outperforms the other methods, demonstrating strong generalization across all evaluated datasets.
Figure 5: Effect of varying inference timesteps $t_{\text{test}}$ across multiple datasets. Each plot shows the top-1 accuracy trend on the NTU-60 and NTU-120 datasets under different splits. The solid red line represents the average accuracy of our method, with the shaded orange area indicating the variation in accuracy across 10 different random Gaussian noise instances. Dashed blue line corresponds to the second-best method in each benchmark.
4.4 Ablation Studies
Effect of loss function design. The ablation study shown in Table 3 evaluates the impact of combining the diffusion loss $\mathcal{L}_{\text{diff}}$ and the triplet diffusion (TD) loss $\mathcal{L}_{\text{TD}}$ on the model’s performance. The results demonstrate that leveraging both losses yields superior performance compared to using either one alone. Specifically, when only $\mathcal{L}_{\text{diff}}$ is employed, the model focuses on accurately denoising the skeleton features conditioned on the text prompts but may lack sufficient discriminative power between similar actions. Conversely, applying only $\mathcal{L}_{\text{TD}}$ enhances discriminative fusion but without ensuring optimal noise prediction. The combination of both losses strikes a balance, ensuring discriminative fusion, which resulting in the highest performance across all evaluated splits.
  $\mathcal{L}_{\text{diff}}$ $\mathcal{L}_{\text{TD}}$ NTU-60 (Acc, %) NTU-120  (Acc, %)
55/5 split 48/12 split 110/10 split 96/24 split
79.87 53.03 72.44 57.65
80.90 54.36 70.73 60.95
86.49 56.03 74.15 65.06
Table 3: Top-1 accuracy of different loss function configurations on the NTU-60 and NTU-120 datasets. The results compare models trained with only the diffusion loss $\mathcal{L}_{\text{diff}}$, only the triplet diffusion loss $\mathcal{L}_{\text{TD}}$, and the combination of both losses.
Impact of total timesteps $T$. The results of the ablation study on diffusion timesteps are shown in Table 4. We observe that the choice of $T$ has a significant impact on performance across all datasets and splits. When $T$ is too small, the model tends to overfit, as the problem becomes too simple, limiting the diversity in noise added to the skeleton features during training. On the other hand, too large $T$ values introduces diverse noise strengths, making it challenging for the model to denoise effectively, which deteriorates performance. The best $T$ is found empirically with $T=50$, striking a balance between maintaining a challenging task and avoiding overfitting.
Impact of timestep $t_{\text{test}}$ and noise $\bm{\epsilon}_{\text{test}}$ in inference. We conducted experiments across different test timesteps $t_{\text{test}}\in[0,50]$ and observed the accuracy trends on multiple datasets, as illustrated in Fig. 5. Based on these observations, we set $t_{\text{test}}=25$ for all experiments. To examine the impact of noise during inference, we repeated experiments with 10 different random Gaussian noise samples. In Fig. 5, the shaded orange regions in the graphs depict the variances in top-1 accuracy due to noise, while the red lines represent the average accuracy. The blue dashed lines indicate the second-best method’s accuracy for comparison. Our analysis shows that while noise variations can cause up to a ±$\pm 2.5\%$-point changes in top-1 accuracy at $t_{\text{test}}=25$, our TDSM shows consistently outperforming the state-of-the-art methods regardless of noise levels.
  Total $T$ NTU-60 (Acc, %) NTU-120  (Acc, %)
55/5 split 48/12 split 110/10 split 96/24 split
1 85.03 44.10 69.91 60.35
10 84.51 50.89 69.97 62.04
50 86.49 56.03 74.15 65.06
100 83.48 56.27 71.05 64.57
500 81.34 53.43 71.93 60.81
Table 4: Top-1 accuracy with varying total timesteps $T$ on the NTU-60 and NTU-120 datasets.
4.5 Limitations
Sensitivity to noise variation. As illustrated in Fig. 5, although our method achieves superior performance, its results exhibit some sensitivity to noise $\bm{\epsilon}_{\text{test}}$ during inference. Recent studies [17] have suggested that predicting the initial state $\mathbf{z}_{x}$ during the reverse diffusion process yields more stable results compared to direct noise ($\bm{\epsilon}$) prediction, especially under varying noise conditions. As part of future work, we plan to explore this refinement to enhance the robustness against noise fluctuations.
5 Conclusion
Our TDSM is the first framework to apply diffusion models to zero-shot skeleton-based action recognition. The selective denoising process promotes a robust fusion of skeleton and text features, allowing the model to develop a discriminative feature space that can generalize to unseen action labels. Also, our approach enhances discriminative fusion through the TD loss which is designed to denoise GT skeleton-text pairs effectively while preventing the fusion of incorrect pairs within the seen dataset. Extensive experiments show that our TDSM significantly outperforms the very recent SOTA models with large margins for various benchmark datasets.